OpenStudy (anonymous):

1. What is the value of 4P2? 6 8 12 24

OpenStudy (anonymous):

Use: \[\large ^nP_r = \frac{n!}{(n-r)!}\]

OpenStudy (anonymous):

Here n = 4 and r = 2..

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

What is \(4!\) ??

OpenStudy (anonymous):

i cant figure it out

OpenStudy (anonymous):

You will get; \[\large ^4P_2 = \frac{4!}{2!} \implies \frac{4 \times 3 \times 2!}{2!}\] After cancelling what are you left with?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Good..

OpenStudy (anonymous):

so its 12?

OpenStudy (anonymous):

how do i do this one > 10C6

OpenStudy (anonymous):

\[\large ^nC_r = \frac{n!}{(n-r)! \cdot r!}\]

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

Work more..

OpenStudy (anonymous):

151,200?

OpenStudy (anonymous):

Now work less.. :)